range of function f(x)= 6+3x+6-x+3-x+2?
a.[-2,infinity)  b.(-2,infinity)  c.(6,infinity)  d.[6,infinity)

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,

f(x) ={ 6^x + 6^-x } + { 3^x + 3^-x } + 2

              Range of { 6^x + 6^-x} = [2,infinity)

              Range of { 3^x + 3^-x } = [2, infinity)

        Hence min value of f(x) = 2+2+2 = 6

           Range = [6, infinity ) .

 

 

Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)